OpenStudy (anonymous):

ok since i am always answering questions i might as well ask one... express in simplest form.

OpenStudy (anonymous):

\[(\sqrt{3}+4)(\sqrt{3}-4\]

OpenStudy (anonymous):

omg i don't know who to tag...

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@Luigi0210

OpenStudy (asnaseer):

do you recognise the formula for the difference between two squares here?

OpenStudy (anonymous):

@Nnesha

OpenStudy (asnaseer):

i.e. \(a^2-b^2=(a+b)(a-b)\)

OpenStudy (anonymous):

nope.... it was something about \(a^2+ab+b^2=(a+b)(a-b)\) ?

OpenStudy (anonymous):

wait was that the equation?

OpenStudy (asnaseer):

not quite:\[(a+b)(a-b)=a(a-b)+b(a-b)=a^2-ab+ab-b^2=a^2-b^2\]

OpenStudy (anonymous):

ok thank you... so does that mean it's \(3-16\)?

OpenStudy (asnaseer):

perfect! :)

OpenStudy (anonymous):

ok.... thanks~~~

OpenStudy (asnaseer):

yw :)

OpenStudy (anonymous):

:D